5. Find the sum:   (b) (37) + (–2) + (–65) + (–8)
hkmfgthjnsrxhge

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(37) + (-2) + (-65) + (-8) = (+37) + (-75)

= (+37) + (-37) + (-38)
= 0 + (-38) [(-a) + (+a) = 0] = -38



Thanks